OpenStudy (anonymous):

How do you change x(t) = sin x and y(t) = cos t from parametric form to non-parametric xy form?

OpenStudy (amistre64):

inverse of x is equal to t plug into y in place of t

OpenStudy (anonymous):

so it would be y = cos (sin^-1 x)?

OpenStudy (amistre64):

i like that

OpenStudy (anonymous):

I'm really not understanding. Can you explain a little plz

OpenStudy (anonymous):

do you recognize \((cos(t),\sin(t))\) as points on the unit circle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know the equation for the unit circle?

OpenStudy (anonymous):

\[x^2+y^2=1\]

OpenStudy (anonymous):

yup

OpenStudy (amistre64):

(sint, cost) is the setup here tho, prolly doesnt make a difference tho

OpenStudy (anonymous):

So the answer is simply the equation of a circle wit a radius of 1

OpenStudy (amistre64):

as long as we are not trying to define a function .... yes

OpenStudy (anonymous):

okay thanks